# Cognitio Emergens: Agency, Dimensions, and Dynamics in Human–AI Knowledge Co-Creation

### v1

@@ 3.4.4 Organizational Implications
The framework underscores how organizational contexts significantly shape human-AI epistemic partnerships, addressing gaps identified in the literature (Liu & Shen, 2025; Lou et al., 2025). Factors such as structural configurations (e.g., modularity, boundary permeability (Joseph & Sengul, 2024)), coordination mechanisms, incentive systems, and cultural elements (e.g., risk tolerance, interpretive diversity) critically influence how partnerships develop, which dimensions flourish, and what they ultimately achieve.
This suggests that optimizing human-AI collaboration requires attention not just to technological capabilities and individual interactions, but crucially to the surrounding organizational design and culture. Cultivating principles of organizational ambidexterity (O’Reilly & Tushman, 2013)—balancing exploration and exploitation, fostering psychological safety for experimentation, and establishing appropriate governance—becomes essential for realizing the potential of advanced agency configurations and dimensional capabilities.
@@ 3.4.5 Ethical Implications
Finally, the framework surfaces profound ethical questions. The Axiological Intelligence dimension suggests that scientific values themselves can evolve through human-AI interaction, potentially creating new axiological frameworks that transcend current ethical categories. This raises questions about how we guide this evolution responsibly.
The tension between innovative potential (especially in Partnership Agency) and the risk of Epistemic Alienation highlights a central ethical challenge: How do we balance the pursuit of transformative discovery with the need to maintain meaningful human understanding, interpretation, and accountability in knowledge creation (Seeber et al., 2018)? This question becomes increasingly urgent as AI systems evolve from assistants to co-authors (Beel, Kan, & Baumgart, 2025; Yamada et al., 2025), requiring us to reconsider the epistemic and ethical foundations of scientific practice.
@@ 3.5 Conclusion
The Cognitio Emergens framework offers a comprehensive theoretical structure for understanding the emergent properties of human-AI epistemic partnerships. By identifying Agency Configurations, Epistemic Dimensions, and Partnership Dynamics, the framework provides conceptual tools for analyzing current partnerships while guiding their future development.
Unlike frameworks focused primarily on technical capabilities or ethical constraints, Cognitio Emergens explicitly addresses the emergent, co-creative potential of human-AI collaboration in knowledge production. It acknowledges both unprecedented opportunities for epistemic transformation and significant challenges that may arise as these partnerships evolve.
The framework’s distinctive integration of visual representation with conceptual structure creates a powerful tool for both theoretical exploration and practical application. The capability signatures approach provides a systematic method for understanding the uneven, sometimes surprising ways that human-AI research partnerships develop across dimensions and agency configurations.
As AI systems continue to evolve from instrumental tools to epistemic partners, frameworks like Cognitio Emergens become increasingly essential for understanding and guiding this transformation. By providing a balanced perspective that neither uncritically celebrates nor unnecessarily fears advanced AI in knowledge production, the framework offers a foundation for developing human-AI partnerships that maximize transformative potential while maintaining epistemic integrity.
The true power of the framework lies not in classification but in transformation—providing a map for navigating the landscape of possibility in human-AI knowledge creation. It reveals the hidden topology of research evolution, showing paths toward breakthrough capabilities that might otherwise remain undiscovered in the complex terrain of emerging human-AI partnerships.
@@ 4. Implementing the Cognitio Emergens Framework
This section bridges theoretical concepts with practical application, translating the Cognitio Emergens framework into actionable implementation guidance. We address three critical implementation domains: configuring agency relationships (4.1), cultivating epistemic dimensions (4.2), and managing partnership dynamics (4.3). For each domain, we provide specific actions and their corresponding value, offering a flexible foundation that research teams can adapt to their unique contexts. While Appendix A provides more comprehensive strategies and assessment protocols, these core principles offer an essential starting point for organizations seeking to implement the framework.
@@ 4.1 Configuring Agency Relationships
Successfully implementing agency configurations—the foundation of the CE framework—requires deliberate design choices that strategically balance innovation potential with interpretive control across different research contexts. Successfully navigating these configurations allows teams to leverage AI appropriately for different research phases and goals.
@@
  • Actions: Define explicit constraints for AI tasks (e.g., permitted data sources, banned methodologies, required output formats). Develop systematic verification protocols aligned with disciplinary standards (e.g., checklists for result validation, independent replication procedures).
  • Value: This ensures AI contributions enhance efficiency without compromising established epistemic rigor, maintaining clear human authority.
@@
  • Actions: Schedule dedicated exploratory sessions where unprompted AI insights (e.g., anomaly detection reports, alternative hypothesis suggestions) can be considered based on potential rather than immediate fit. Develop evaluation criteria that balance novelty with epistemic grounding (e.g., rating AI suggestions on plausibility and originality). Implement workflows where AI suggestions trigger specific human review processes.
  • Value: This allows teams to benefit from AI’s potential to spark innovation and challenge assumptions, moving beyond single-loop learning without sacrificing critical evaluation.
@@
  • Actions: Establish rapid, iterative co-creation workflows where human and AI contributions build upon each other with minimal delay (e.g., using shared digital whiteboards or collaborative coding environments). Foster interpretive communities—groups of researchers who collectively examine, debate, and interpret partnership outputs—to build shared ownership and navigate complexity. Develop shared conceptual vocabularies or ‘partnership dialects’ to facilitate communication about emergent concepts.
  • Value: This configuration maximizes the potential for breakthrough insights emerging from the synergy of human and AI cognition, enabling triple-loop learning and paradigm shifts.
@@
  • Actions: Identify early signals suggesting configuration shifts (e.g., increasing interpretive challenges signaling a need to retreat; breakthrough AI suggestions signaling opportunities to advance). Develop team capacity (e.g., through training and reflective practice) to operate across multiple configurations, consciously selecting the mode best suited to the current task (e.g., using Partnership for brainstorming, Directed for validation).
@@
While configuring appropriate agency relationships establishes the basic structural conditions for productive collaboration, cultivating specific epistemic dimensions develops the partnership’s functional capabilities. These dimensions represent the specific collaborative capacities that emerge through sustained interaction.
@@ 4.2 Cultivating Epistemic Dimensions
Developing the partnership’s functional capabilities requires deliberate cultivation of the six epistemic dimensions—specific collaborative capacities that emerge through sustained interaction and targeted development efforts. This involves fostering specific practices and environments that encourage the development of desired capabilities, leading to more potent and balanced partnerships.
@@
  • Divergent Intelligence: Foster by creating psychologically safe environments supporting novel hypothesis generation (e.g., brainstorming sessions where AI outputs are initially treated as provocations, not solutions). Implement variation techniques (e.g., prompt engineering that explicitly asks AI for outlier perspectives or counter-arguments) to expand the space of considered possibilities. Value: Enhances the partnership’s ability to generate potentially groundbreaking ideas.
  • Interpretive Intelligence: Develop protocols for generating different types of explanations calibrated to specific research needs (e.g., requesting simplified summaries for quick checks, detailed causal trace explanations for critical validation). Utilize interactive visualization tools that allow researchers to explore AI models or outputs. Implement regular ‘interpretability audits’ where the team assesses the understandability of recent AI contributions. Value: Builds trust, facilitates deeper collaboration, and crucially mitigates the risk of epistemic alienation.
@@
  • Connective Intelligence: Implement cross-domain exposure practices (e.g., tasking AI to explicitly search for analogous problems or solutions in adjacent fields; dedicating team time to exploring AI-identified cross-disciplinary links). Utilize network visualization tools to map connections surfaced by AI across disparate literatures or datasets. Value: Facilitates interdisciplinary breakthroughs and avoids reinventing wheels.
  • Synthesis Intelligence: Organize integration workshops where teams explicitly work to combine diverse perspectives (human theoretical insights, AI data patterns) into coherent explanatory frameworks. Use multi-level modeling approaches capable of incorporating findings at different scales or from different methodologies. Task AI with generating draft syntheses or conceptual maps based on provided inputs. Value: Develops more comprehensive and robust understandings by integrating fragmented knowledge.
@@
  • Anticipatory Intelligence: Develop structured scenario exploration approaches (e.g., using AI to generate multiple plausible future trajectories based on different assumptions; conducting ‘pre-mortem’ analyses on potential research pathways). Implement ‘early warning systems’ where AI monitors trends for indicators of potential disruptions or emerging opportunities. Value: Enables more proactive and strategic research planning, especially in fast-moving fields.
  • Axiological Intelligence: Conduct structured dialogues examining the epistemic values embedded in research choices and AI outputs (e.g., using AI to surface implicit assumptions in assessment criteria; team discussions on what constitutes a ‘significant’ finding). Co-develop new evaluation criteria with AI that reflect the unique contributions of the partnership. Value: Fosters critical reflection on research goals and ensures alignment between partnership activities and deeper scientific values, potentially reshaping disciplinary norms.
Implementation approaches should target dimensions strategically based on the partnership’s current capability profile and goals, rather than attempting uniform advancement. Focusing on dimensions that complement existing strengths or address critical weaknesses allows teams to develop balanced capability signatures tailored to their specific research contexts, maximizing their collaborative potential.
@@ 4.3 Managing Partnership Dynamics
  • Actions: Implement practices that periodically challenge disciplinary paradigms (e.g., ‘what if’ sessions guided by AI provocations). Explicitly document the partnership’s evolution, including shifts in understanding, methodology, or interaction protocols, fostering Recursive Evolution. Dedicate resources to exploring high-risk, high-reward avenues suggested by the partnership.
@@
  • Actions: Establish practices for modulating research pace based on epistemic needs (e.g., scheduled ‘slow thinking’ sessions for deep interpretation; defined ‘sprints’ for rapid exploration). Implement clear protocols for switching between exploratory modes (prioritizing novelty) and validation modes (prioritizing rigor). Foster Epistemic Ambidexterity through regular reflection on the balance between innovation and interpretability.
@@
  • Actions: Develop practices for detecting early signs of Epistemic Alienation (e.g., researcher reluctance to explain AI outputs, divergence between formal endorsement and intuitive understanding) and intervene (e.g., through focused interpretability efforts or temporary shifts to more Directed agency). Mitigate Epistemic Closure by actively seeking diverse inputs (e.g., using multiple AI models, involving external reviewers, designing AI prompts to challenge consensus) and periodically auditing the partnership for reinforcing biases.
@@ 4.4 Conclusion
By thoughtfully applying and refining these approaches, teams can cultivate collaborative relationships that harness AI’s transformative potential while preserving meaningful human participation and epistemic integrity.
@@ 5. Discussion
The Cognitio Emergens framework represents a significant advancement in understanding human-AI epistemic partnerships, offering both theoretical insight and practical guidance for navigating this emerging landscape. While previous sections detailed the framework’s components and implementation, this discussion synthesizes its broader significance, revealing how CE fundamentally transforms our understanding of knowledge creation while providing actionable guidance for stakeholders. We explore the framework’s integrative theoretical contributions, its practical implications for different audiences, promising directions for future inquiry, and important considerations regarding its scope and boundaries.
@@ 5.1 Integrative Theoretical Significance
While section 3.4 detailed specific theoretical implications, here we examine how these implications collectively constitute a paradigm shift in understanding human-AI collaboration. The Cognitio Emergens framework offers five integrative insights that fundamentally transform our understanding of knowledge creation in the age of advanced AI: boundary transcendence, recursive structure-emergence relationships, constitutive temporal evolution, theoretical integration, and bridging descriptive-normative perspectives. Together, these insights reframe how we conceptualize, study, and guide human-AI epistemic partnerships.
5.1.1 Transcending Traditional Boundaries
The framework’s most profound contribution—boundary transcendence—lies in moving beyond rigid distinctions between human knowers and AI tools to reveal a hybrid cognitive system where knowledge emerges from structural coupling and interaction. In essence, it describes a hybrid cognitive system. This boundary-crossing manifests simultaneously across epistemology (new forms of distributed knowing), agency (blurring lines of control and contribution), and methodology (integrating diverse approaches).
This quality connects CE to broader intellectual movements challenging dualistic thinking (e.g., ecological approaches to cognition, organizational theories of permeable boundaries). Just as these movements reconfigured understanding in their domains, CE reconfigures our understanding of scientific knowledge creation beyond the individual human researcher or purely social processes.
@@ 5.1.2 Recursive Relationship Between Structure and Emergence
The framework reveals a fundamental recursivity in human-AI partnerships—a mutually influential relationship between structural configurations (Agency) and emergent capabilities (Dimensions, Dynamics) that challenges linear models of collaboration. Agency configurations enable capabilities; these capabilities reshape agency; dynamics influence both. This self-reinforcing cycle challenges linear, mechanistic models of human-AI collaboration.
This recursive pattern connects CE to complexity theory and studies of emergence. Human-AI partnerships function as complex adaptive systems. They generate higher-order properties through component interactions. Importantly, these interactions create feedback loops that transform the components themselves. This means guiding these partnerships requires attending to both structural conditions and emergent properties simultaneously, rather than focusing solely on component design or outcome metrics.
@@ 5.1.3 Temporal Evolution as Constitutive Rather than Incidental
CE positions evolution as fundamental to human-AI relationships, not merely an add-on feature. Agency configurations oscillate, dimensions develop unevenly, and dynamics create evolving cycles. Time and change are inherent.
This temporal orientation connects CE to evolutionary approaches, but with a key difference: CE emphasizes how partnerships actively construct the epistemic environments they inhabit through evolving practices, rather than just adapting to fixed external conditions. This implies that sustainable, productive partnerships require deliberate cultivation of developmental trajectories—balancing adaptation with intentional shaping of the collaborative process over time.

### v2

@@ 3.4.4 Organizational Implications
The framework underscores how organizational contexts shape human-AI epistemic partnerships, addressing gaps identified in the literature (Lou et al. 2025; Liu and Shen 2025). Factors such as structural configurations (e.g., modularity, boundary permeability (Joseph and Sengul 2024)), coordination mechanisms, incentive systems, and cultural elements (e.g., risk tolerance, interpretive diversity) influence how partnerships develop, which dimensions flourish, and what they achieve.
This suggests that optimizing human-AI collaboration requires attention beyond technological capabilities and individual interactions to surrounding organizational design and culture. Cultivating principles of organizational ambidexterity (O’Reilly and Tushman 2013)—balancing exploration and exploitation, fostering psychological safety for experimentation, and establishing appropriate governance—becomes essential for realizing potential of advanced agency configurations and dimensional capabilities.
@@ 3.4.5 Ethical Implications
The framework surfaces ethical questions. The Axiological Intelligence dimension suggests that scientific values themselves can evolve through human-AI interaction, potentially creating new axiological frameworks that transcend current ethical categories. This raises questions about how we guide this evolution responsibly.
The tension between innovative potential (especially in Partnership Agency) and risk of Epistemic Alienation highlights a central ethical challenge: How do we balance pursuit of transformative discovery with need to maintain meaningful human understanding, interpretation, and accountability in knowledge creation (Seeber et al. 2018)? This question becomes increasingly urgent as AI systems evolve from assistants to co-authors (Beel, Kan, and Baumgart 2025; Yamada et al. 2025), requiring us to reconsider epistemic and ethical foundations of scientific practice.
@@ 3.5 Conclusion
The Cognitio Emergens framework offers a theoretical structure for understanding emergent properties of human-AI epistemic partnerships. By identifying Agency Configurations, Epistemic Dimensions, and Partnership Dynamics, the framework provides conceptual tools for analyzing current partnerships while guiding their future development.
Unlike frameworks focused primarily on technical capabilities or ethical constraints, Cognitio Emergens explicitly addresses the emergent, co-creative potential of human-AI collaboration in knowledge production. It acknowledges both opportunities for epistemic transformation and challenges that may arise as partnerships evolve.
The framework’s integration of visual representation with conceptual structure creates a tool for both theoretical exploration and practical application. The capability signatures approach provides a systematic method for understanding the uneven ways that partnerships develop across dimensions and agency configurations.
As AI systems evolve from instrumental tools to epistemic partners, frameworks like Cognitio Emergens become essential for understanding and guiding this transformation. The framework offers a foundation for developing partnerships that maximize transformative potential while maintaining epistemic integrity.
@@ 4. Implementing the Cognitio Emergens Framework
This section bridges theoretical concepts with practical application, translating the Cognitio Emergens framework into actionable implementation guidance. We address three implementation domains: configuring agency relationships (4.1), cultivating epistemic dimensions (4.2), and managing partnership dynamics (4.3). For each domain, we provide specific actions and their corresponding value, offering a flexible foundation that research teams can adapt to their unique contexts. While Appendix A provides more comprehensive strategies and assessment protocols, these core principles offer a starting point for organizations seeking to implement the framework.
@@ 4.1 Configuring Agency Relationships
Implementing agency configurations—the foundation of the CE framework—requires deliberate design choices that balance innovation potential with interpretive control across different research contexts. Navigating these configurations allows teams to leverage AI appropriately for different research phases and goals.
@@
  • Actions: Define explicit constraints for AI tasks (permitted data sources, banned methodologies, required output formats). Develop systematic verification protocols aligned with disciplinary standards (checklists for result validation, independent replication procedures).
  • Value: Ensures AI contributions enhance efficiency without compromising established epistemic rigor, maintaining clear human authority.
@@
  • Actions: Schedule dedicated exploratory sessions where unprompted AI insights (anomaly detection reports, alternative hypothesis suggestions) can be considered based on potential rather than immediate fit. Develop evaluation criteria that balance novelty with epistemic grounding (rating AI suggestions on plausibility and originality). Implement workflows where AI suggestions trigger specific human review processes.
  • Value: Allows teams to benefit from AI’s potential to spark innovation and challenge assumptions, moving beyond single-loop learning without sacrificing critical evaluation.
@@
  • Actions: Establish rapid, iterative co-creation workflows where human and AI contributions build upon each other with minimal delay (using shared digital whiteboards or collaborative coding environments). Foster interpretive communities—groups of researchers who collectively examine, debate, and interpret partnership outputs—to build shared ownership and navigate complexity. Develop shared conceptual vocabularies or ‘partnership dialects’ to facilitate communication about emergent concepts.
  • Value: Maximizes potential for breakthrough insights emerging from the synergy of human and AI cognition, enabling triple-loop learning and paradigm shifts.
@@
  • Actions: Identify early signals suggesting configuration shifts (increasing interpretive challenges signaling a need to retreat; breakthrough AI suggestions signaling opportunities to advance). Develop team capacity (through training and reflective practice) to operate across multiple configurations, consciously selecting the mode best suited to the current task (using Partnership for brainstorming, Directed for validation).
@@
While configuring appropriate agency relationships establishes structural conditions for productive collaboration, cultivating specific epistemic dimensions develops the partnership’s functional capabilities. These dimensions represent collaborative capacities that emerge through sustained interaction.
@@ 4.2 Cultivating Epistemic Dimensions
Developing the partnership’s functional capabilities requires deliberate cultivation of the six epistemic dimensions—collaborative capacities that emerge through sustained interaction and targeted development efforts. This involves fostering specific practices and environments that encourage development of desired capabilities, leading to more potent and balanced partnerships.
@@
  • Divergent Intelligence: Foster by creating psychologically safe environments supporting novel hypothesis generation (brainstorming sessions where AI outputs are initially treated as provocations, not solutions). Implement variation techniques (prompt engineering that explicitly asks AI for outlier perspectives or counter-arguments) to expand the space of considered possibilities. Value: Enhances ability to generate potentially groundbreaking ideas.
  • Interpretive Intelligence: Develop protocols for generating different types of explanations calibrated to specific research needs (requesting simplified summaries for quick checks, detailed causal trace explanations for critical validation). Utilize interactive visualization tools that allow researchers to explore AI models or outputs. Implement regular ‘interpretability audits’ where the team assesses understandability of recent AI contributions. Value: Builds trust, facilitates deeper collaboration, and mitigates the risk of epistemic alienation.
@@
  • Connective Intelligence: Implement cross-domain exposure practices (tasking AI to explicitly search for analogous problems or solutions in adjacent fields; dedicating team time to exploring AI-identified cross-disciplinary links). Utilize network visualization tools to map connections surfaced by AI across disparate literatures or datasets. Value: Facilitates interdisciplinary breakthroughs and avoids reinventing wheels.
  • Synthesis Intelligence: Organize integration workshops where teams explicitly work to combine diverse perspectives (human theoretical insights, AI data patterns) into coherent explanatory frameworks. Use multi-level modeling approaches capable of incorporating findings at different scales or from different methodologies. Task AI with generating draft syntheses or conceptual maps based on provided inputs. Value: Develops more comprehensive understandings by integrating fragmented knowledge.
@@
  • Anticipatory Intelligence: Develop structured scenario exploration approaches (using AI to generate multiple plausible future trajectories based on different assumptions; conducting ‘pre-mortem’ analyses on potential research pathways). Implement ‘early warning systems’ where AI monitors trends for indicators of potential disruptions or emerging opportunities. Value: Enables more proactive and strategic research planning, especially in fast-moving fields.
  • Axiological Intelligence: Conduct structured dialogues examining the epistemic values embedded in research choices and AI outputs (using AI to surface implicit assumptions in assessment criteria; team discussions on what constitutes a ‘significant’ finding). Co-develop new evaluation criteria with AI that reflect the unique contributions of the partnership. Value: Fosters critical reflection on research goals and ensures alignment between partnership activities and deeper scientific values, potentially reshaping disciplinary norms.
Implementation approaches should target dimensions strategically based on the partnership’s current capability profile and goals, rather than attempting uniform advancement. Focusing on dimensions that complement existing strengths or address weaknesses allows teams to develop balanced capability signatures tailored to their specific research contexts, maximizing their collaborative potential.
@@ 4.3 Managing Partnership Dynamics
  • Actions: Implement practices that periodically challenge disciplinary paradigms (‘what if’ sessions guided by AI provocations). Explicitly document the partnership’s evolution, including shifts in understanding, methodology, or interaction protocols, fostering Recursive Evolution. Dedicate resources to exploring high-risk, high-reward avenues suggested by the partnership.
@@
  • Actions: Establish practices for modulating research pace based on epistemic needs (scheduled ‘slow thinking’ sessions for deep interpretation; defined ‘sprints’ for rapid exploration). Implement clear protocols for switching between exploratory modes (prioritizing novelty) and validation modes (prioritizing rigor). Foster Epistemic Ambidexterity through regular reflection on the balance between innovation and interpretability.
@@
  • Actions: Develop practices for detecting early signs of Epistemic Alienation (researcher reluctance to explain AI outputs, divergence between formal endorsement and intuitive understanding) and intervene (through focused interpretability efforts or temporary shifts to more Directed agency). Mitigate Epistemic Closure by actively seeking diverse inputs (using multiple AI models, involving external reviewers, designing AI prompts to challenge consensus) and periodically auditing the partnership for reinforcing biases.
@@ 5. Discussion
The Cognitio Emergens framework offers theoretical insight and practical guidance for navigating human-AI epistemic partnerships. While previous sections detailed the framework’s components and implementation, this discussion synthesizes its broader significance. We explore the framework’s integrative theoretical contributions, practical implications for different audiences, directions for future inquiry, and considerations regarding its scope and boundaries.
@@ 5.1 Integrative Theoretical Significance
While section 3.4 detailed specific theoretical implications, here we examine how these implications collectively shift understanding of human-AI collaboration. The Cognitio Emergens framework offers five integrative insights: boundary transcendence, recursive structure-emergence relationships, constitutive temporal evolution, theoretical integration, and bridging descriptive-normative perspectives. Together, these insights reframe how we conceptualize, study, and guide human-AI epistemic partnerships.
5.1.1 Beyond Traditional Boundaries
Boundary transcendence moves beyond rigid distinctions between human knowers and AI tools to reveal a hybrid cognitive system where knowledge emerges from structural coupling and interaction. This boundary-crossing manifests simultaneously across epistemology (new forms of distributed knowing), agency (blurring lines of control and contribution), and methodology (integrating diverse approaches).
This connects CE to broader intellectual movements challenging dualistic thinking (ecological approaches to cognition, organizational theories of permeable boundaries). Just as these movements reconfigured understanding in their domains, CE reconfigures understanding of scientific knowledge creation beyond the individual human researcher or purely social processes.
@@ 5.1.2 Recursive Relationship Between Structure and Emergence
Human-AI partnerships exhibit recursivity: structural configurations (Agency) and emergent capabilities (Dimensions, Dynamics) mutually influence each other in ways that challenge linear models of collaboration. Agency configurations enable capabilities; these capabilities then reshape agency; dynamics influence both simultaneously.
This recursive pattern connects CE to complexity theory and studies of emergence. Human-AI partnerships function as complex adaptive systems. They generate higher-order properties through component interactions. These interactions create feedback loops that transform the components themselves. Guiding these partnerships requires attending to both structural conditions and emergent properties simultaneously, rather than focusing solely on component design or outcome metrics.
@@ 5.1.3 Temporal Evolution as Constitutive Rather than Incidental
CE positions evolution as central to human-AI relationships. Agency configurations oscillate, dimensions develop unevenly, and dynamics create evolving cycles. Time and change are inherent.
This temporal orientation connects CE to evolutionary approaches, with a key difference: CE emphasizes how partnerships actively construct the epistemic environments they inhabit through evolving practices, rather than just adapting to fixed external conditions. This implies that sustainable partnerships require deliberate cultivation of developmental trajectories—balancing adaptation with intentional shaping of the collaborative process over time.